### v2

# The chromatic number of 4-dimensional lattices

Abstract.
The chromatic number of a lattice in $n$-dimensional Euclidean space is defined as the chromatic number of its Voronoi graph. The Voronoi graph is the Cayley graph on the lattice having the strict Voronoi vectors as generators. In this paper we determine the chromatic number of all $4$-dimensional lattices.
To achieve this we use the known classification of $52$ parallelohedra in dimension $4$. These $52$ geometric types yield $16$ combinatorial types of relevant Voronoi graphs. We discuss a systematic approach to checking for isomorphism of Cayley graphs of lattices.
Lower bounds for the chromatic number are obtained from choosing appropriate small finite induced subgraphs of the Voronoi graphs. Matching upper bounds are derived from periodic colorings. To determine the chromatic numbers of these finite graphs, we employ a SAT solver.
Dedicated to Rudolf Scharlau in occasion of his 72nd birthday
1. Introduction
Let $\Lambda\subseteq\mathbb{R}^{n}$ be an $n$-dimensional lattice in $n$-dimensional Euclidean space. The chromatic number of $\Lambda$ is the smallest number of colors one needs to color lattice translates of the Voronoi cell
$V(\Lambda)=\{x\in\mathbb{R}^{n}:\|x\|\leq\|x-v\|\text{ for all }v\in\Lambda\}$
so that distinct lattice translates $v+V(\Lambda)$ and $w+V(\Lambda)$, with $v\neq w$, receive different colors whenever the intersection $(v+V(\Lambda))\cap(w+V(\Lambda))$ is a common $(n-1)$-dimensional face (a facet) of the two polytopes $v+V(\Lambda)$ and $w+V(\Lambda)$. This chromatic number is denoted by $\chi(\Lambda)$.
The chromatic number of lattices was first considered by Dutour Sikirić, Madore, Moustrou, Vallentin in [8].
Equivalently, the chromatic number of a lattice $\Lambda$ is the chromatic number of the (infinite) Cayley graph on the additive group $\Lambda$ with strict Voronoi vectors as its set of generators. The strict Voronoi vectors of $\Lambda$, denoted by $\operatorname{Vor}(\Lambda)$, are the differences $v-w$ such that the intersection of the distinct translates $v+V(\Lambda)$ and $w+V(\Lambda)$ is a common facet. Thus,
$\chi(\Lambda)=\chi(\operatorname{Cayley}(\Lambda,\operatorname{Vor}(\Lambda))),$
where the vertices of the Cayley graph are all elements of $\Lambda$ and two distinct vertices $v,w$ are adjacent whenever the difference $v-w$ lies in $\operatorname{Vor}(\Lambda)$. We call the Cayley graph $\operatorname{Cayley}(\Lambda,\operatorname{Vor}(\Lambda))$ the Voronoi graph of the lattice $\Lambda$.
Many natural questions regarding $\chi(\Lambda)$ remain unanswered. For instance, it is not known whether $\chi(\Lambda)$ is computable. By the compactness theorem of de Bruijn and Erdős [3] we only know that the decision problem “Is $\chi(\Lambda)\leq k$?” is semidecidable [^1] There is an algorithm which terminates with a positive answer after finitely many steps on every input $\Lambda$ with $\chi(\Lambda)\leq k$. On inputs $\Lambda$ with $\chi(\Lambda)>k$, the algorithm either terminates with a negative answer or it runs in an infinite loop. because $\chi(\Lambda)$ is equal to the largest chromatic number of all finite, induced subgraphs of $\Lambda$. Directly related is the question of whether one can always find a periodic coloring with $\chi(\Lambda)$ colors. If so, and if we knew a priori a bound on the period, then this would immediately yield an algorithm for computing $\chi(\Lambda)$.
Also the extremal question of maximizing $\chi(\Lambda)$ among all $n$-dimensional lattices $\Lambda$ is widely open. In [8] it is shown, using a combination of the Kabatiansky-Levenshtein upper bound and the Minkowski-Hlawka lower bound for sphere packings, that this maximum grows exponentially for generic lattices. In particular, there are $n$-dimensional lattices with $\chi(\Lambda_{n})\geq 2\cdot 2^{(0.0990\ldots-o(1))n}$. By an easy greedy argument, there is also an exponential upper bound $\chi(\Lambda)\leq 2^{n}$ for every $n$-dimensional lattice $\Lambda$. Currently, an efficient way (polynomial time in the dimension) to construct lattices with exponential chromatic number is not known.
Turning to low dimensions, in [8] the two- and three-dimensional cases are settled completely. There are two cases in dimension two: the square lattice and the hexagonal lattice, with chromatic number two and three, respectively. There are five cases in dimension three. The Voronoi cells of these five cases are the cube, the hexagonal prism, the rhombic dodecahedron, the elongated dodecahedron, and the truncated octahedron. The corresponding chromatic numbers are two, three, and four (three times).
In this paper we completely settle the four-dimensional case.
Theorem 1.1.
If $\Lambda$ is a four-dimensional lattice, then its Voronoi graph is isomorphic to one of the 16 cases given in Table 2.4.222After the first version of our paper was submitted to the arXiv.org e-print archive, Igor Baburin contacted us and reported that he arrived at the same number a few years ago. The 16 pairwise non-isomorphic graphs have chromatic numbers as given in Table 3.1.
The proof consists of two parts.
In Section 2 we show that, up to graph isomorphism, there are $16$ different Voronoi graphs to consider. This classification is a consequence of a classical result in the geometry of numbers, essentially due to Delaunay [5]: In dimension $4$, there are $52$ types of Voronoi cells. We prove that these $52$ types only yield $16$ non-isomorphic Voronoi graphs. In Section 2 we describe these classifications in detail. In addition, we present an argument that shows how to algorithmically test the isomorphism of Cayley graphs of lattices. Since these are infinite graphs, it is not a priori clear that such an algorithm needs to exist.
In Section 3 we consider each of the $16$ cases separately. For every case we first determine lower bounds for $\chi(\Lambda)$ by computing the chromatic number of a finite, induced subgraph of the Voronoi graph $\operatorname{Cayley}(\Lambda,\operatorname{Vor}(\Lambda))$. Every natural number $d$ defines a finite, induced subgraph of the Voronoi graph by using the vertices which are at most (graph) distance $d$ from the origin. We denote this graph by $\operatorname{Cayley}(\Lambda,\operatorname{Vor}(\Lambda))_{d}$. Then, for each case, we determine matching upper bounds by constructing explicit periodic colorings. The computations for the lower bounds and for the upper bounds both use SAT solvers—algorithms to certify whether a given Boolean formula in conjunctive normal form has a satisfying assignment or not.
The proof of Theorem 1.1 involves numerous computations. To ensure transparency and enable the reader to verify our calculations, we have provided a program, written using the computer algebra system MAGMA [2]. This MAGMA program is available as an ancillary file from the arXiv.org e-print archive. We discuss the MAGMA program at the end of Section 3 once all the needed objects have been introduced.
In Section 4 we collect some questions for further research.
2. Voronoi graphs of lattices in four dimensions
In this section we show that there are, up to graph isomorphism, $16$ Voronoi graphs $\operatorname{Cayley}(\Lambda,\operatorname{Vor}(\Lambda))$ where $\Lambda$ is a four-dimensional lattice.
The starting point for this classification of Voronoi graphs is a classical classification result by Voronoi [17] concerning the different Delaunay subdivisions of $4$ dimensional lattices.
Note that there is another possible starting point with the classification of iso-edge domains (also called $C$-type domains). For more information on the concept of iso-edge domains we refer to Ryshkov and Baranovskii [13] and to Dutour Sikirić and Kummer [9]. In general we now have $3$ equivalence relations regarding lattices, by identifying lattices according to either their Voronoi graph, their iso-edge domain, or their Delaunay subdivision. It is a priori clear that these relations become finer with Delaunay subdivisions being the finest one and Voroni graphs being the coarsest one. In dimension $4$ we have, up to equivalence, $51$ inequivalent iso-edge domains and $52$ inequivalent Delaunay subdivisions, so for our approach there is not much of a difference in using one of the two. We prefer to work with Delaunay subdivisions, as they are more common in the literature. Note that our classification of Voroni graphs in dimension $4$ yields that there are $16$ such, up to equivalence, so their number is strictly smaller then the number of iso-edge domains.
We now return to the discussion of Delaunay subdivisions. To state Voronoi’s result it is convenient to work with positive definite quadratic forms instead of lattices.
By $\mathcal{S}^{n}_{++}$ we denote the open set of positive definite quadratic forms, which we identify by positive definite matrices through $x\mapsto Q[x]=x^{\sf T}Qx$. The dictionary between lattices in Euclidean space and positive definite quadratic forms can be compactly expressed by the double coset $\operatorname{O}_{n}(\mathbb{R})\backslash\operatorname{GL}_{n}(\mathbb{R})/\operatorname{GL}_{n}(\mathbb{Z})$, representing Euclidean lattices up to orthogonal transformations and lattice basis transformations, which is equal to $\mathcal{S}^{n}_{++}/\operatorname{GL}_{n}(\mathbb{Z})$, positive definite quadratic forms up to the conjugation action by $\operatorname{GL}_{n}(\mathbb{Z})$, as $\operatorname{O}_{n}(\mathbb{R})\backslash\operatorname{GL}_{n}(\mathbb{R})=\mathcal{S}^{n}_{++}$.
Below we define the Delaunay subdivision of a positive definite quadratic form. This polytopal subdivision is geometrically dual to the subdivision of $\mathbb{R}^{n}$ given by lattice translates of Voronoi cells.
A polytope $P\subseteq\mathbb{R}^{n}$ with only integral vertices is called a Delaunay polytope of $Q\in\mathcal{S}^{n}_{++}$ if there exists $c\in\mathbb{R}^{n}$ and $r>0$ such that
$Q[x-c]\geq r\text{ for all $x\in\mathbb{Z}^{n}$ }\;\text{ and }\;Q[x-c]=r\text{ for all vertices $x$ of $P$}.$
The set
$\operatorname{Del}(Q)=\{P:\text{$P$ is a Delaunay polytope of $Q$}\}$
gives the Delaunay subdivision of $Q$. If all occurring Delaunay polytopes are simplices we speak of a Delaunay triangulation.
Now we shall explain the connection between Delaunay subdivisions and Voronoi graphs of lattices. Let $\Lambda$ be a lattice with lattice basis $B\in\operatorname{GL}_{n}(\mathbb{R})$. When we take the linear image, under $B$, of the vertex-edge graph of the Delaunay subdivision of $Q=B^{\sf T}B$, then we exactly get the Voronoi graph $\operatorname{Cayley}(\Lambda,\operatorname{Vor}(\Lambda))$. So classifying Delaunay subdivisions up to the action of $\operatorname{GL}_{n}(\mathbb{Z})$ provides a finer classification than the classification of Voronoi graphs up to the graph isomorphism.
In his second memoir [17], Voronoi showed that up to the action of $\operatorname{GL}_{4}(\mathbb{Z})$ there are exactly three different Delaunay triangulations of quaternary positive definite quadratic forms. We recall some details of his result in Section 2.1.
Every Delaunay subdivision $\mathcal{D}$ is refined by some Delaunay triangulation $\mathcal{D}^{\prime}$, meaning that all elements of $\mathcal{D}^{\prime}$ are simplices and for every simplex $D^{\prime}\in\mathcal{D}^{\prime}$ there is a polytope $D\in\mathcal{D}$ so that $D^{\prime}\subseteq D$. So classifying Delaunay subdivisions amounts to finding all possible Delaunay subdivisions which coarsen some Delaunay triangulation. The classification of Delaunay subdivisions of quaternary positive definite quadratic forms is due to Delaunay [5]. Interestingly, he used the dual approach of classifying four-dimensional parallelohedra.
Parallelohedra are polytopes which tile space by lattice translates. For dimension $4$, Delaunay showed that parallelohedra are affine images of Voronoi cells of lattices. In general, it is not known whether every parallelohedron arises as an affine image of a Voronoi cell. This question is called Voronoi’s conjecture.
Delaunay claimed that there are $51$ different Delaunay subdivisions of quaternary positive definite quadratic forms. However, there are indeed $52$ cases; Delaunay’s classification was corrected and completed by Shtogrin [15]. Later Delaunay’s result was reworked, verified, and simplified by Engel [10], Conway [4], Deza, Grishukhin [6], Zhilinskii [18]. We will recall some details of Conway’s classification in Section 2.2. More details, and a complete derivation of Conway’s classification, can be found in the first-named author’s PhD thesis [16].
In Section 2.3 we will show that many of the $52$ cases yield the same or at least isomorphic Voronoi graphs.
2.1. Classification of Delaunay triangulations
Here we recall some details of Voronoi’s classification of Delaunay triangulations of quaternary positive definite quadratic forms. We refer to the book by Schürmann [14] for a contemporary exhibition of Voronoi’s theory.
Let $\mathcal{D}$ be the Delaunay subdivision of some positive definite quadratic form $Q$, then
$\Delta(\mathcal{D})=\{Q^{\prime}\in\mathcal{S}^{n}_{++}:\mathcal{D}=\operatorname{Del}(Q^{\prime})\}$
is called the secondary cone of $\mathcal{D}$. Its topological closure $\overline{\Delta(\mathcal{D})}$ is a polyhedral cone in the convex cone of positive semidefinite matrices $\mathcal{S}^{n}_{+}$. The Delaunay subdivision is a triangulation if and only if its secondary cone is full-dimensional. A Delaunay subdivision $\mathcal{D}^{\prime}$ refines another subdivision $\mathcal{D}$ if and only if $\overline{\Delta(\mathcal{D^{\prime}})}\subseteq\overline{\Delta(\mathcal{D})}$. The group $\operatorname{GL}_{n}(\mathbb{Z})$ acts on the space of symmetric matrices by conjugation. Under this group action there are only finitely many orbits of secondary cones.
At the end of his second memoir, Voronoi [17] computed the orbits of full-dimensional secondary cones for $n\leq 4$. For $n\leq 3$ there is only one orbit and for $n=4$ there are exactly three orbits.
Using Voronoi’s notation, the three polyhedral domains ${\it\Delta}$, ${\it\Delta^{\prime}}$, ${\it\Delta^{\prime\prime}}$ are given by the extremal rays:
$\begin{array}[]{c}{\it\Delta}=\operatorname{cone}\{R_{1},\ldots,R_{10}\},\\ {\it\Delta^{\prime}}=\operatorname{cone}\{R_{1},\ldots,R_{4},R_{6},\ldots,R_{11}\},\\ {\it\Delta^{\prime\prime}}=\operatorname{cone}\{R_{1},\ldots,R_{4},R_{6},\ldots,R_{9},R_{11},R_{12}\},\end{array}$
where
$R_{1}=\begin{pmatrix}1&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix},\;R_{2}=\begin{pmatrix}0&0&0&0\\ 0&1&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix},R_{3}=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&1&0\\ 0&0&0&0\end{pmatrix},$
$R_{4}=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&1\end{pmatrix},\;R_{5}=\begin{pmatrix}1&-1&0&0\\ -1&1&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix},\;R_{6}=\begin{pmatrix}1&0&-1&0\\ 0&0&0&0\\ -1&0&1&0\\ 0&0&0&0\end{pmatrix},$
$R_{7}=\begin{pmatrix}1&0&0&-1\\ 0&0&0&0\\ 0&0&0&0\\ -1&0&0&1\end{pmatrix},\;R_{8}=\begin{pmatrix}0&0&0&0\\ 0&1&-1&0\\ 0&-1&1&0\\ 0&0&0&0\end{pmatrix},\;R_{9}=\begin{pmatrix}0&0&0&0\\ 0&1&0&-1\\ 0&0&0&0\\ 0&-1&0&1\end{pmatrix},$
$R_{10}=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&1&-1\\ 0&0&-1&1\end{pmatrix},\;R_{11}=\begin{pmatrix}4&2&-2&-2\\ 2&4&-2&-2\\ -2&-2&4&0\\ -2&-2&0&4\end{pmatrix},\;R_{12}=\begin{pmatrix}1&1&-1&-1\\ 1&1&-1&-1\\ -1&-1&1&1\\ -1&-1&1&1\end{pmatrix}.$
In our notation this corresponds to
${\it\Delta}=\overline{\Delta(\mathcal{D})},\;{\it\Delta^{\prime}}=\overline{\Delta(\mathcal{D}^{\prime})},\;{\it\Delta^{\prime\prime}}=\overline{\Delta(\mathcal{D}^{\prime\prime})},$
where $\mathcal{D}$, $\mathcal{D}^{\prime}$, $\mathcal{D}^{\prime\prime}$ are non-equivalent Delaunay triangulations. In Figure 2.1 we schematically show how the rational closure of the positive definite quadratic forms is tessellated by orbits of the polyhedral domains ${\it\Delta}$, ${\it\Delta^{\prime}}$, ${\it\Delta^{\prime\prime}}$.
Figure 2.1. A finite part of the tessellation of the rational closure of the positive definite quadratic forms into orbits of the polyhedral domains ${\it\Delta}$, ${\it\Delta^{\prime}}$, ${\it\Delta^{\prime\prime}}$. Black nodes correspond to the orbit of ${\it\Delta}$, grey nodes to the orbit of ${\it\Delta^{\prime}}$, and white nodes to the orbit of ${\it\Delta^{\prime\prime}}$. We draw an edge between two nodes whenever the polyhedral domains share a facet, here a face of dimension $9$.
2.2. Classification of Delaunay subdivisions
There are $52$ different Delaunay subdivisions of quaternary positive definite quadratic forms. This classification can be derived from Voronoi’s result. To accomplish this, one considers all faces of the three simplicial cones ${\it\Delta}$, ${\it\Delta^{\prime}}$, ${\it\Delta^{\prime\prime}}$ which gives $\leq 3072$ candidates, represented by $\sum_{i=1}^{12}\alpha_{i}R_{i}$, with $\alpha_{i}\in\{0,1\}$. Then, one checks which of these faces contain positive definite quadratic forms in their relative interiors and which ones are equivalent under the action of $\operatorname{GL}_{4}(\mathbb{Z})$.
In Table 2.2 and Table 2.3, we give the classification symbols of Conway. For every representative positive definite quadratic form, we require the strict Voronoi vectors, which we list in Table 2.1. For more details on this classification and an explanation of the symbols used we refer to [4] and [16].
$\begin{split}&v_{1}=\pm(0,0,0,1),\;v_{2}=\pm(0,0,1,-1),\;v_{3}=\pm(0,0,1,0),\;v_{4}=\pm(0,0,1,1),\\ &v_{5}=\pm(0,1,0,0),\;v_{6}=\pm(0,1,0,1),\;v_{7}=\pm(0,1,1,0),\;v_{8}=\pm(0,1,1,1),\\ &v_{9}=\pm(1,-1,0,0),\;v_{10}=\pm(1,0,0,0),\;v_{11}=\pm(1,0,0,1),\;v_{12}=\pm(1,0,1,0),\\ &v_{13}=\pm(1,0,1,1),\;v_{14}=\pm(1,1,0,0),\;v_{15}=\pm(1,1,0,1),\;v_{16}=\pm(1,1,1,0),\\ &v_{17}=\pm(1,1,1,1)\end{split}$
Table 2.1. The complete list of all possible strict Voronoi vectors of positive definite quadratic forms in ${\it\Delta}$, ${\it\Delta^{\prime}}$, ${\it\Delta^{\prime\prime}}$.
lattice dimension representative
number strict Voronoi vectors
$K_{5}$ $10$ $R_{1}+R_{2}+R_{3}+R_{4}+R_{5}+R_{6}+R_{7}+R_{8}+R_{9}+R_{10}$
$30$ $v_{1},v_{3},v_{4},v_{5},v_{6},v_{7},v_{8},v_{10},v_{11},v_{12},v_{13},v_{14},v_{15},v_{16},v_{17}$
$K_{3,3}$ $9$ $R_{1}+R_{2}+R_{3}+R_{4}+R_{6}+R_{7}+R_{8}+R_{9}+R_{12}$
$30$ $v_{1},v_{2},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{15},v_{16},v_{17}$
$K_{5}-1$ $9$ $R_{1}+R_{2}+R_{3}+R_{4}+R_{5}+R_{7}+R_{8}+R_{9}+R_{10}$
$28$ $v_{1},v_{3},v_{4},v_{5},v_{6},v_{7},v_{8},v_{10},v_{11},v_{13},v_{14},v_{15},v_{16},v_{17}$
$K_{5}-2$ $8$ $R_{1}+R_{2}+R_{3}+R_{4}+R_{7}+R_{8}+R_{9}+R_{10}$
$24$ $v_{1},v_{3},v_{4},v_{5},v_{6},v_{7},v_{8},v_{10},v_{11},v_{13},v_{15},v_{17}$
$K_{5}-1-1$ $8$ $R_{1}+R_{2}+R_{3}+R_{4}+R_{5}+R_{7}+R_{8}+R_{10}$
$26$ $v_{1},v_{3},v_{4},v_{5},v_{7},v_{8},v_{10},v_{11},v_{13},v_{14},v_{15},v_{16},v_{17}$
$K_{5}-3$ $7$ $R_{1}+R_{2}+R_{4}+R_{7}+R_{8}+R_{9}+R_{10}$
$20$ $v_{1},v_{3},v_{4},v_{5},v_{7},v_{8},v_{10},v_{11},v_{13},v_{17}$
$K_{5}-2-1$ $7$ $R_{1}+R_{2}+R_{4}+R_{5}+R_{7}+R_{8}+R_{10}$
$24$ $v_{1},v_{3},v_{4},v_{5},v_{7},v_{8},v_{10},v_{11},v_{13},v_{14},v_{16},v_{17}$
$K_{4}+1$ $7$ $R_{1}+R_{2}+R_{3}+R_{4}+R_{8}+R_{9}+R_{10}$
$16$ $v_{1},v_{3},v_{4},v_{5},v_{6},v_{7},v_{8},v_{10}$
$C_{2221}$ $7$ $R_{1}+R_{2}+R_{3}+R_{4}+R_{7}+R_{9}+R_{10}$
$22$ $v_{1},v_{3},v_{4},v_{5},v_{6},v_{8},v_{10},v_{11},v_{13},v_{15},v_{17}$
$C_{221}+1$ $6$ $R_{1}+R_{2}+R_{3}+R_{4}+R_{8}+R_{10}$
$14$ $v_{1},v_{3},v_{4},v_{5},v_{7},v_{8},v_{10}$
$C_{321}$ $6$ $R_{1}+R_{2}+R_{4}+R_{7}+R_{8}+R_{10}$
$20$ $v_{1},v_{3},v_{4},v_{5},v_{7},v_{8},v_{10},v_{11},v_{13},v_{17}$
$C_{222}$ $6$ $R_{1}+R_{2}+R_{3}+R_{7}+R_{9}+R_{10}$
$22$ $v_{1},v_{3},v_{4},v_{5},v_{6},v_{8},v_{10},v_{11},v_{13},v_{15},v_{17}$
$C_{3}+C_{3}$ $6$ $R_{1}+R_{4}+R_{7}+R_{8}+R_{9}+R_{10}$
$12$ $v_{3},v_{5},v_{7},v_{8},v_{10},v_{17}$
$C_{5}$ $5$ $R_{1}+R_{2}+R_{7}+R_{8}+R_{10}$
$20$ $v_{1},v_{3},v_{4},v_{5},v_{7},v_{8},v_{10},v_{11},v_{13},v_{17}$
$C_{4}+1$ $5$ $R_{1}+R_{2}+R_{4}+R_{8}+R_{10}$
$14$ $v_{1},v_{3},v_{4},v_{5},v_{7},v_{8},v_{10}$
$C_{3}+1+1$ $5$ $R_{1}+R_{2}+R_{3}+R_{4}+R_{8}$
$10$ $v_{1},v_{3},v_{5},v_{7},v_{10}$
$1+1+1+1$ $4$ $R_{1}+R_{2}+R_{3}+R_{4}$
$8$ $v_{1},v_{3},v_{5},v_{10}$
Table 2.2. Representatives of the $17$ different four-dimensional Delaunay subdivisions where the corresponding Voronoi cell is a zonotope, that is, a linear projection of a cube. In Conway’s classification the graph $K_{4}$ occurs, which is wrong and should be replaced by $C_{221}+1$; this mistake was found by Deza and Grishukhin in [6]. The entry “dimension” gives the dimension of the secondary cone of the corresponding Delaunay subdivision. The entry “number” gives the number of strict Voronoi vectors.
lattice dimension representative
number strict Voronoi vectors
$111+$ $10$ $R_{1}+R_{2}+R_{3}+R_{4}+R_{6}+R_{7}+R_{8}+R_{9}+R_{11}+R_{12}$
$30$ $v_{1},v_{2},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{15},v_{16},v_{17}$
$111-$ $10$ $R_{1}+R_{2}+R_{3}+R_{4}+R_{6}+R_{7}+R_{8}+R_{9}+R_{10}+R_{11}$
$30$ $v_{1},v_{3},v_{4},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{15},v_{16},v_{17}$
$211+$ $9$ $R_{1}+R_{2}+R_{3}+R_{4}+R_{6}+R_{8}+R_{9}+R_{11}+R_{12}$
$28$ $v_{1},v_{2},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{16},v_{17}$
$211-$ $9$ $R_{1}+R_{2}+R_{3}+R_{4}+R_{6}+R_{8}+R_{9}+R_{10}+R_{11}$
$28$ $v_{1},v_{3},v_{4},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{16},v_{17}$
$311+$ $8$ $R_{1}+R_{2}+R_{3}+R_{4}+R_{6}+R_{8}+R_{11}+R_{12}$
$28$ $v_{1},v_{2},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{16},v_{17}$
$311-$ $8$ $R_{1}+R_{2}+R_{3}+R_{4}+R_{6}+R_{8}+R_{10}+R_{11}$
$28$ $v_{1},v_{3},v_{4},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{16},v_{17}$
$221+$ $8$ $R_{1}+R_{2}+R_{3}+R_{4}+R_{8}+R_{9}+R_{11}+R_{12}$
$26$ $v_{1},v_{2},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
$221-$ $8$ $R_{1}+R_{2}+R_{3}+R_{4}+R_{8}+R_{9}+R_{10}+R_{11}$
$26$ $v_{1},v_{3},v_{4},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
$122^{\prime}$ 1 $8$ $R_{1}+R_{2}+R_{3}+R_{4}+R_{7}+R_{8}+R_{11}+R_{12}$
$26$ $v_{1},v_{2},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17},$
$411$ $7$ $R_{1}+R_{2}+R_{4}+R_{6}+R_{8}+R_{10}+R_{11}$
$28$ $v_{1},v_{3},v_{4},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{16},v_{17}$
$321+$ $7$ $R_{1}+R_{2}+R_{4}+R_{7}+R_{8}+R_{10}+R_{11}$
$26$ $v_{1},v_{3},v_{4},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
$321-$ $7$ $R_{1}+R_{2}+R_{3}+R_{4}+R_{8}+R_{10}+R_{11}$
$26$ $v_{1},v_{3},v_{4},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
$222+$ $7$ $R_{1}+R_{3}+R_{4}+R_{8}+R_{9}+R_{11}+R_{12}$
$24$ $v_{1},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
$222-$ $7$ $R_{1}+R_{3}+R_{4}+R_{6}+R_{7}+R_{10}+R_{11}$
$24$ $v_{1},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
$222^{\prime}$ $7$ $R_{1}+R_{3}+R_{4}+R_{8}+R_{9}+R_{10}+R_{11}$
$24$ $v_{1},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
$22^{\prime}2^{\prime\prime}$ $7$ $R_{1}+R_{4}+R_{7}+R_{8}+R_{9}+R_{10}+R_{11}$
$24$ $v_{1},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
Table 2.3. (First part) Representatives of the $35$ different four-dimensional Delaunay subdivisions where the corresponding Voronoi cell is not a zonotope.
lattice dimension representative
number strict Voronoi vectors
$421$ $6$ $R_{1}+R_{2}+R_{4}+R_{8}+R_{10}+R_{11}$
$26$ $v_{1},v_{3},v_{4},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
$331+$ $6$ $R_{1}+R_{2}+R_{3}+R_{4}+R_{11}+R_{12}$
$26$ $v_{1},v_{2},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
$331-$ $6$ $R_{1}+R_{2}+R_{3}+R_{4}+R_{10}+R_{11}$
$26$ $v_{1},v_{3},v_{4},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17},$
$322+$ $6$ $R_{1}+R_{3}+R_{4}+R_{8}+R_{11}+R_{12}$
$24$ $v_{1},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
$322-$ $6$ $R_{1}+R_{4}+R_{7}+R_{8}+R_{11}+R_{12}$
$24$ $v_{1},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
$322^{\prime}$ $6$ $R_{1}+R_{2}+R_{4}+R_{6}+R_{7}+R_{11}$
$24$ $v_{1},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
$431$ $5$ $R_{1}+R_{2}+R_{4}+R_{10}+R_{11}$
$26$ $v_{1},v_{3},v_{4},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
$422$ $5$ $R_{1}+R_{4}+R_{8}+R_{11}+R_{12}$
$24$ $v_{1},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
$422^{\prime}$ $5$ $R_{1}+R_{4}+R_{8}+R_{10}+R_{11}$
$24$ $v_{1},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
$332+$ $5$ $R_{1}+R_{3}+R_{4}+R_{11}+R_{12}$
$24$ $v_{1},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
$332-$ $5$ $R_{1}+R_{3}+R_{4}+R_{10}+R_{11}$
$24$ $v_{1},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
$441$ $5$ $R_{1}+R_{2}+R_{10}+R_{11}$
$26$ $v_{1},v_{3},v_{4},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
$432$ $4$ $R_{1}+R_{4}+R_{10}+R_{11}$
$24$ $v_{1},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
$333+$ $4$ $R_{3}+R_{4}+R_{11}+R_{12}$
$24$ $v_{1},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
$333-$ $4$ $R_{3}+R_{4}+R_{10}+R_{11}$
$24$ $v_{1},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
$442$ $3$ $R_{1}+R_{10}+R_{11}$
$24$ $v_{1},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
$433$ $3$ $R_{4}+R_{10}+R_{11}$
$24$ $v_{1},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
$443$ $2$ $R_{10}+R_{11}$
$24$ $v_{1},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
$444$ $1$ $R_{11}$
$24$ $v_{1},v_{3},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10},v_{11},v_{12},v_{13},v_{17}$
Table 2.3. (Second part) Representatives of the $35$ different four-dimensional Delaunay subdivisions where the corresponding Voronoi cell is not a zonotope.
2.3. Classification of Voronoi graphs
In this section, we derive a classification of Voronoi graphs for four-dimensional lattices based on the classification of Delaunay subdivisions. We have to verify which of the above $52$ Voronoi graphs are isomorphic.
In many cases, the strict Voronoi vectors coincide, so that also the Voronoi graphs coincide. In three cases, for example, for the pairs $(411,311+)$, $(441,331+)$, $(K_{5}-2,K_{5}-2-1)$ the strict Voronoi vectors do not coincide, but the Voronoi graphs are still isomorphic. In the first two cases one can find an isomorphism by determining an element $A\in\operatorname{GL}_{4}(\mathbb{Z})$ of the automorphism group of $444$ (which corresponds to the root lattice $D_{4}$ and which has order $1152$) so that the Voronoi vectors of $411$ are mapped to the ones of $311+$, respectively of $441$ to $331+$. This linear isomorphism between the generators of the Cayley graph induces a global graph isomorphism. In the third case, one can use an element of the automorphism group of $K_{5}$ (which corresponds to the weight lattice $A^{*}_{4}$ and which has order $240$).
Then we are left with $16$ candidates of pairwise non-isomorphic Voronoi graphs. They are indeed pairwise non-isomorphic, which one can see by looking at invariants. The first invariant we are using is the regularity $r$ of the graphs. The second and third invariant we are using are coming from the finite, induced subgraph $\operatorname{Cayley}(\Lambda,\operatorname{Vor}(\Lambda))_{1}$ whose vertices are at most unit distance from the origin. The second invariant is the number of edges $|E|$ of $\operatorname{Cayley}(\Lambda,\operatorname{Vor}(\Lambda))_{1}$ and the third invariant is the order of the (graph) automorphism group of $\operatorname{Cayley}(\Lambda,\operatorname{Vor}(\Lambda))_{1}$.
# graph $r$ $|E|$ order number representatives
1 $V_{30}^{z}$ $30$ $180$ $240$ $1$ $K_{5}$
2 $V_{30}^{z,n}$ $30$ $186$ $144$ $2$ $K_{3,3}$, $111+$
3 $V_{30}^{n}$ $30$ $180$ $24$ $1$ $111-$
4 $V_{28}^{z}$ $28$ $154$ $24$ $1$ $K_{5}-1$
5 $V_{28}^{n}$ $28$ $160$ $24$ $5$ $211+$, $211-$, $311+$, $311-$, $411$
6 $V_{26}^{z}$ $26$ $134$ $16$ $1$ $K_{5}-1-1$
7 $V_{26}^{n}$ $26$ $140$ $96$ $10$ $221+$, $221-$, $122^{\prime}$ 1, $321+$, $321-$,
$421$, $331+$, $331-$, $431$, $441$
8 $V_{24}^{z}$ $24$ $114$ $16$ $2$ $K_{5}-2$, $K_{5}-2-1$
9 $V_{24}^{n}$ $24$ $120$ $1152$ $18$ $222+$, $222-$, $222^{\prime}$, $22^{\prime}2^{\prime\prime}$, $322+$,
$322-$, $322^{\prime}$, $422$, $422^{\prime}$, $332+$,
$332-$, $432$, $333+$, $333-$, $442$,
$433$, $443$, $444$
10 $V_{22}^{z}$ $22$ $94$ $96$ $2$ $C_{2221}$, $C_{222}$
11 $V_{20}^{z}$ $20$ $80$ $240$ $3$ $K_{5}-3$, $C_{321}$, $C_{5}$
12 $V_{16}^{z}$ $16$ $52$ $96$ $1$ $K_{4}+1$
13 $V_{14}^{z}$ $14$ $38$ $96$ $2$ $C_{221}+1$, $C_{4}+1$
14 $V_{12}^{z}$ $12$ $24$ $288$ $1$ $C_{3}+C_{3}$
15 $V_{10}^{z}$ $10$ $16$ $288$ $1$ $C_{3}+1+1$
16 $V_{8}^{z}$ $8$ $8$ $40320$ $1$ $1+1+1+1$
Table 2.4. Classification of Voronoi graphs of four-dimensional lattices. The column $r$ gives the regularity of the Voronoi graph, the column $|E|$ gives the number of edges, the column “order” gives the order of the automorphism group of the induced subgraph $\operatorname{Cayley}(\Lambda,\operatorname{Vor}(\Lambda))_{1}$.
Here, we succeeded testing whether two Voronoi graphs are isomorphic in a rather ad hoc manner. This can also be done in a systematic way: In the next section we provide a finite algorithm that can be used in general, for this we will briefly discuss how to test for isomorphism of general Cayley graphs associated with lattices with finite sets of generators; see also the paper [1] by Baburin which summarizes computational approaches to finding isomorphisms and automorphisms of Cayley graphs of lattices and contains references to earlier works in this field.
2.4. Isomorphisms of Cayley graphs of lattices
Suppose $\Lambda$ and $\Lambda^{\prime}$ are lattices, and $S\subseteq\Lambda$ and $S^{\prime}\subseteq\Lambda^{\prime}$ are finite, centrally symmetric generating sets. That is, $S=-S$ and $\Lambda=\langle S\rangle_{\mathbb{Z}}$, and the corresponding identities hold for $\Lambda^{\prime}$ and $S^{\prime}$.
Any linear isomorphism between $\Lambda$ and $\Lambda^{\prime}$ that maps $S$ to $S^{\prime}$ induces a graph isomorphism between $\operatorname{Cayley}(\Lambda,S)$ and $\operatorname{Cayley}(\Lambda^{\prime},S^{\prime})$. A general result on isomorphisms of Cayley graphs of finitely generated abelian groups (see Löh [12]) implies that the converse is also true: Any graph isomorphism $\varphi:\operatorname{Cayley}(\Lambda,S)\rightarrow\operatorname{Cayley}(\Lambda^{\prime},S^{\prime})$ induces a linear isomorphism of $\Lambda$ and $\Lambda^{\prime}$ which sends $S$ to $S^{\prime}$.
Since we think this result is of independent interest in the context of lattices, and the elegant proof of the general result can be significantly shortened in this context, we reproduce a shortened version of the proof given in Löh [12] here.
The main idea of the proof is to use induction on the size of the generating set $S$, the crucial observation is that this induction should be done by reducing the size of $S$ by removing an $\|\cdot\|_{2}$-maximal element from $S$.
So, suppose $\Lambda$ is a lattice with a centrally symmetric generating set $S$. Given $s\in S$ we write $S_{s}=S\setminus\{s,-s\}$ and $\Lambda_{s}=\langle S_{s}\rangle_{\mathbb{Z}}$.
We rest the induction step on the following lemma.
Lemma 2.1.
Let $s\in S$ be $\|\cdot\|_{2}$-maximal and let $\varphi:\operatorname{Cayley}(\Lambda,S)\rightarrow\operatorname{Cayley}(\Lambda^{\prime},S^{\prime})$ be a graph isomorphism. Then the restriction $\varphi_{s}$ of $\varphi$ to $\operatorname{Cayley}(\Lambda_{s},S_{s})$ induces a graph isomorphism $\operatorname{Cayley}(\Lambda_{s},S_{s})\rightarrow\operatorname{Cayley}(\Lambda^{\prime}_{s^{\prime}},S^{\prime}_{s^{\prime}})$, where $s^{\prime}=\varphi(s)$.
Proof.
The result will be an immediate consequence of Lemma 2.2
The workhorse of the proof is the following combination of Proposition 2.5 and Proposition 2.10 in [12]. For this we need to introduce some concepts from the theory of (Cayley) graphs. Firstly, an algebraic line of type $s\in S$ in $\operatorname{Cayley}(\Lambda,S)$ is a $\mathbb{Z}$-path $\gamma:\mathbb{Z}\rightarrow\Lambda$ with $n\mapsto v+ns$, for some $v\in\Lambda$. Secondly, a geodesic segment in a graph $\Gamma=(V,E)$ is a finite path $(v_{0},\ldots,v_{n})$ with shortest path distance $d_{\Gamma}(v_{0},v_{n})=n$ and a geodesic line is a $\mathbb{Z}$-path $\gamma:\mathbb{Z}\rightarrow V$, such that $d_{\Gamma}(\gamma(i),\gamma(j))=|i-j|$ for all $i,j\in\mathbb{Z}$. Lastly, a convex geodesic line in $\Gamma$ is a geodesic line $\gamma$, such that for all $m,n\in\mathbb{Z}$ with $m\leq n$ there is a unique geodesic segment in $\Gamma$ starting at $\gamma(m)$ and ending at $\gamma(n)$.
Lemma 2.2.
Let $\Lambda,\Lambda^{\prime}$ be lattices and $S,S^{\prime}$ be finite, centrally symmetric generating sets of $\Lambda,\Lambda^{\prime}$ and let $\varphi:\operatorname{Cayley}(\Lambda,S)\rightarrow\operatorname{Cayley}(\Lambda^{\prime},S^{\prime})$ be a graph isomorphism. Then:
  1. (1)
    If $s$ is a $\|\cdot\|_{2}$-maximal element of $S$, then all algebraic lines of type $s$ are convex geodesic lines.
  2. (2)
    $\varphi$ maps convex geodesic lines to convex geodesic lines.
  3. (3)
    Every convex geodesic line is algebraic.
  4. (4)
    If $\gamma$ and $\eta$ are algebraic lines of type $s$ in $\operatorname{Cayley}(\Lambda,S)$, then $\varphi\circ\gamma$ and $\varphi\circ\eta$ are algebraic lines of the same type in $\operatorname{Cayley}(\Lambda^{\prime},S^{\prime})$. In particular, their type is $s^{\prime}=\varphi(s)$.
Proof.
(1) through (3) make up Proposition 2.5 in [12].
We give a short sketch of the proof of (1) to highlight the importance of $\|\cdot\|_{2}$-maximality of $s$.
For this choose an arbitrary geodesic segment $(g_{0},\ldots,g_{k})$ in $\operatorname{Cayley}(\Lambda,S)$ connecting points $\gamma(m),\gamma(n)$ on $\gamma$ (assume $m\leq n$). Write
$(n-m)s=\gamma(n)-\gamma(m)=\sum_{i=0}^{k-1}g_{i+1}-g_{i}$
where $g_{i+1}-g_{i}\in S$. Then by the maximality of $s$ we can conclude, using the triangle inequality, that $n-m=k$, $\|g_{i+1}-g_{i}\|_{2}=\|s\|_{2}$ and then $g_{i+1}-g_{i}=s$. From this it follows that the arbitrary geodesic segment $(g_{0},\ldots,g_{k})$ is a segment on $\gamma$ as needed.
(4) is a simplification of Proposition 2.10 in [12]. We present a streamlined version of the proof here: Write $\gamma^{\prime}=\varphi\circ\gamma$ and $\eta^{\prime}=\varphi\circ\eta$ and elementwise for $n\in\mathbb{Z}$
$\gamma(n)$ $=v+ns$
$\eta(n)$ $=w+ns$
$\gamma^{\prime}(n)$ $=v^{\prime}+ns^{\prime}$
$\eta^{\prime}(n)$ $=w^{\prime}+nt^{\prime}.$
We have to show $s^{\prime}=t^{\prime}$. We consider the graph distances $d$ and $d^{\prime}$ on $\operatorname{Cayley}(\Lambda,S)$ and $\operatorname{Cayley}(\Lambda^{\prime},S^{\prime})$. Since $\varphi$ is an isometry and the graph distance is translation invariant we can compute
$d^{\prime}(n(s^{\prime}-t^{\prime}),w^{\prime}-v^{\prime})$ $=d^{\prime}(\gamma^{\prime}(n),\eta^{\prime}(n))$
$=d(\gamma(n),\eta(n))$
$=d(n(s-s),w-v)=d(0,w-v)$
and observe that the right hand side is independent of $n$. Therefore, the set $\{n(s^{\prime}-t^{\prime})\ :\ n\in\mathbb{Z}\}$ is contained in a ball of finite radius around $w^{\prime}-v^{\prime}$ and is hence finite. This implies that $s^{\prime}-t^{\prime}$ has finite order in $\Lambda$ and since $\Lambda$ is torsion free, $s^{\prime}-t^{\prime}=0$. ∎
Now we can prove the isomorphism characterization.
Theorem 2.3.
Let $\Lambda,\Lambda^{\prime}$ be lattices and $S,S^{\prime}$ be finite centrally symmetric generating sets of $\Lambda,\Lambda^{\prime}$. If $\varphi:\operatorname{Cayley}(\Lambda,S)\rightarrow\operatorname{Cayley}(\Lambda^{\prime},S^{\prime})$ is a graph isomorphism, then $\varphi$ is an affine map. In particular, if $\varphi(0)=0$, then $\varphi$ is linear.
Proof.
We can assume that $\varphi(0)=0$, since this can always be achieved by composing $\varphi$ with a suitable translation.
The proof then proceeds by induction on the size of $S$. Let $s\in S$ be a $\|\cdot\|_{2}$-maximal element of $S$. By Lemma 2.1 we know that the restriction $\varphi_{s}$ induces a graph isomorphism $\operatorname{Cayley}(\Lambda_{s},S_{s})\cong\operatorname{Cayley}(\Lambda^{\prime}_{s^{\prime}},S^{\prime}_{s^{\prime}})$. By induction hypothesis, this is a linear isomorphism.
Now we show that this implies that also $\varphi$ is linear. For this take arbitrary $v,w\in\Lambda$ and write them as
$v=x+ms,\quad w=y+ns$
with $x,y\in\Lambda_{s}$, $m,n\in\mathbb{Z}$. Then
$\varphi(v)=\varphi(x)+ms^{\prime},\ \varphi(w)=\varphi(y)+ns^{\prime},\ \text{and}\ \varphi(v+w)=\varphi(x+y)+(m+n)s^{\prime}.$
Since $\varphi$ acts on $\Lambda_{s}$ as $\varphi_{s}$, and the latter is additive, we obtain
$\varphi(v+w)=\varphi(x+y)+(m+n)s^{\prime}=(\varphi(x)+ms^{\prime})+(\varphi(y)+ns^{\prime})=\varphi(v)+\varphi(w).$
So $\varphi$ is indeed additive. ∎
This approach suggests a finite algorithm to test isomorphism of Cayley graphs of two lattices: Let $\Lambda$ and $\Lambda^{\prime}$ be lattices with finite centrally symmetric generating sets $S$ and $S^{\prime}$. For simplicity, assume that both $\Lambda$ and $\Lambda^{\prime}$ are full-dimensional in their respective dimensions.
  1. (1)
    First, fix an arbitrary (vector space) basis contained in $S$, say $s_{1},\ldots,s_{n}$.
  2. (2)
    Second, consider all linear maps defined by sending $s_{1},\ldots,s_{n}$ to $s^{\prime}_{1},\ldots,s^{\prime}_{n}$, where the latter runs through all bases contained in $S^{\prime}$.
  3. (3)
    For each such map $\varphi$ we now check whether $S$ is mapped onto $S^{\prime}$. If that is the case, $\varphi$ is an isomorphism of the associated Cayley graphs.
This algorithm terminates, because $S^{\prime}$ is finite. It is correct because Theorem 2.3 asserts that any isomorphism of the associated Cayley graphs is of this form.
3. Bounds for $\chi(\Lambda)$ and proof of Theorem 1.1
Now we consider each of the $16$ cases identified in Section 2.3 separately. Assume $C=\operatorname{Cayley}(\Lambda,\operatorname{Vor}(\Lambda))$ is the Cayley graph of a lattice $\Lambda$.
First, in Section 3.1, we will consider the finite induced subgraph
$C_{1}=\operatorname{Cayley}(\Lambda,\operatorname{Vor}(\Lambda))_{1}$
with vertices $\{0\}\cup\operatorname{Vor}(\Lambda)$. This subgraph captures the combinatorial structure, the vertex-edge graph, of all Delaunay polytopes of $\Lambda$ containing the origin as a vertex. The chromatic number of this subgraph provides a lower bound for $\chi(\Lambda)$ which we call the Delaunay polytope bound (DPB). We can explicitly compute this bound by reformulating the chromatic number problem in terms of a Boolean satisfiability problem (SAT) and using a SAT solver.
In Section 3.2 we discuss a method to generate periodic colorings of $C$, which we then use to provide an upper bound on $\chi(\Lambda)$, which we call the discrete torus bound (DTB). The validity of the proposed periodic colorings comes from a quotient-like construction of finite graphs and it is again verified through a reformulation as a SAT problem.
In Section 3.3 we explain the computational proof of the main theorem. We recall how to reformulate the determination of the chromatic number as a SAT formula. We also comment on the accompanying Magma program used to perform the involved computational verifications.
3.1. The Delaunay polytope bound
Consider the induced subgraph $C_{1}$ of the Cayley graph $C=\operatorname{Cayley}(\Lambda,\operatorname{Vor}(\Lambda))$ of a lattice $\Lambda$ with vertex set $V_{1}=\{0\}\cup\operatorname{Vor}(\Lambda)$. This is precisely the vertex-edge graph of the polyhedral complex of all Delaunay polytopes of $\Lambda$ containing the origin. Indeed if $[x,y]$ is an edge of any such Delaunay polytope $D$, then $x-y$ is a relevant vector, since the Voronoi cells $x+V(\Lambda)$ and $y+V(\Lambda)$ meet in a common facet by duality of Voronoi and Delaunay tessellations. This gives us the Delaunay polytope bound.
Lemma 3.1.
With the above
$\chi(\Lambda)=\chi(C)\geq\chi(C_{1}).$
$(0,-1)$$(-1,0)$$(-1,-1)$$(0,1)$$(1,1)$$(1,0)$$(0,0)$$c$
Figure 3.1. A Delaunay polytope (red) of the hexagonal lattice included in the vertex-edge graph of the polyhedral complex of all Delaunay polytopes of $\Lambda$ containing the origin and the Voronoi cells centered at its vertices. The red edges are a finite subgraph of the Voronoi graph of the hexagonal lattice and provide a certificate that $\chi(A_{2})\geq 3$.
3.2. The discrete torus bound
The goal of this subsection is to explicitly construct periodic colorings of a lattice $\Lambda$. For this we construct a finite auxiliary object, which we will refer to as a discrete torus of $\Lambda$: Let $\Lambda^{\prime}\subseteq\Lambda$ be a full dimensional sublattice, this induces a graph, the discrete torus of $\Lambda$ by $\Lambda^{\prime}$, with vertex set $V=\Lambda/\Lambda^{\prime}$ and edge set
$E=\{\{x+\Lambda^{\prime},y+\Lambda^{\prime}\}\ :\ \text{there is }v\in\Lambda^{\prime}\text{ with }v+x-y\in\operatorname{Vor}(\Lambda)\}.$
Lemma 3.2.
Let $\Lambda^{\prime}\subseteq\Lambda$ be a sublattice of $\Lambda$ that does not contain any Voronoi relevant vector of $\Lambda$. Then a coloring of Λ/$\Lambda/\Lambda^{\prime}$ can be $\Lambda^{\prime}$-periodically extended to a coloring of $\Lambda$. If $\chi(\Lambda/\Lambda^{\prime})<\infty$ we have $\chi(\Lambda)\leq\chi(\Lambda/\Lambda^{\prime})$.
Proof.
By assumption $\Lambda^{\prime}$ does not contain any relevant Voronoi vectors of $\Lambda$, this implies that every coset $v+\Lambda^{\prime}$ is an independent set in $\operatorname{Cayley}(\Lambda,\operatorname{Vor}(\Lambda))$. Thus any coloring of Λ/$\Lambda/\Lambda^{\prime}$ can be extended $\Lambda^{\prime}$-periodically as claimed and therefore $\chi(\Lambda/\Lambda^{\prime})\geq\chi(\Lambda)$. ∎
For any choice of a full-dimenssional $\Lambda^{\prime}$ we refer to the bound of the above lemma as the discrete torus bound induced by $\Lambda^{\prime}$.
$(0,0)$$(1,0)$$(0,1)$$(1,1)$$(0,2)$$(2,0)$$(1,2)$$(2,1)$$(2,2)$
Figure 3.2. The discrete torus graph $A_{2}/3A_{2}$ of the hexagonal lattice which has $9$ vertices and $27$ edges. This graph gives a certificate that $\chi(A_{2})\leq 3$
3.3. Proof of the main theorem
The proof of our main theorem ultimately involves a finite list of computations. The general strategy is as follows. We separately consider each of the 16 isomorphism classes of Cayley graphs of four-dimensional lattices. For each such graph $C$, we compute the Delaunay polytope bound, that is the chromatic number of the subgraph $C_{1}$ (see Section 3.1). Next we select a lattice $\Lambda$ representing $C$ and choose a suitable sublattice $\Lambda^{\prime}$ such that the discrete torus bound induced by $\Lambda^{\prime}$ matches the Delaunay polytope bound. It turns out, simply by trial and error, that the sublattices $\Lambda^{\prime}$ can be chosen to be of the form $c\Lambda$ for some $c\in\mathbb{N}$, such that the bounds coincide. The information on which lattice and constant $c$ we used for our computation can be found in Table 3.1.
# graph $\Lambda$ $c$ $\chi$
1 $V_{30}^{z}$ $K_{5}$ 5 5
2 $V_{30}^{z,n}$ $K_{3,3}$ 7 7
3 $V_{30}^{n}$ $111-$ 6 6
4 $V_{28}^{z}$ $K_{5}-1$ 5 5
5 $V_{28}^{n}$ $211+$ 6 6
6 $V_{26}^{z}$ $K_{5}-1-1$ 5 5
7 $V_{26}^{n}$ $221+$ 6 6
8 $V_{24}^{z}$ $K_{5}-2-1$ 5 5
9 $V_{24}^{n}$ $222+$ 4 4
10 $V_{22}^{z}$ $C_{2221}$ 4 4
11 $V_{20}^{z}$ $K_{5}-3$ 5 5
12 $V_{16}^{z}$ $K_{4}+1$ 4 4
13 $V_{14}^{z}$ $C_{221}+1$ 4 4
14 $V_{12}^{z}$ $C_{3}+C_{3}$ 3 3
15 $V_{10}^{z}$ $C_{3}+1+1$ 3 3
16 $V_{8}^{z}$ $1+1+1+1$ 2 2
Table 3.1. Computational data used to compute the chromatic number of the 16 cases of inequivalent Cayley graphs associated to $4$ dimensional lattices: For each graph we specify which lattice $\Lambda$ and which constant $c$ we used to compute the discrete torus bound associated to $\Lambda/c\Lambda$. $\chi$ gives the chromatic number of the graph, as obtained by verifying that the Delaunay polytope bound and discrete torus bound coincide.
We also quickly recall how to compute the chromatic number of a finite graph by transforming the problem into a Boolean satisfiability problem, so that a standard SAT solver can be used.
So, let $G=(V,E)$ be a finite graph. Since a coloring of $G$ is a partition of the vertex set into independent sets (the color classes) we have to model:
  1. (1)
    If $v\in V$, then $v$ has a color.
  2. (2)
    If $\{v,w\}\in E$, then $v$ and $w$ do not have the same color.
Note that in principle these two clauses would allow a vertex $v$ to have multiple colors, we do not care about this, as we can then always extract a proper coloring by choosing one of the colors for each multi-colored vertex.
We look for the smallest number $k$ such that there exists a coloring with $k$ color classes. This can be rephrased as a Boolean formula in conjunctive normal for (CNF) as follows. Write $V=\{v_{1},\ldots,v_{n}\}$; these are the variables. Assume that the colors are given by $1,\ldots,k$. We use a total of $n\cdot k$ variables
$\{x_{i,\ell}\}_{\begin{subarray}{c}i=1,\ldots,n\\ \ell=1,\ldots,k\end{subarray}}$
where the value of $x_{i,\ell}$ is to be interpreted by
$x_{i,\ell}=\begin{cases}1,&\text{vertex $v_{i}$ has color $\ell$}\\ 0,&\text{else.}\end{cases}$
With these variables we formulate a SAT problem with two types of clauses:
  1. (1)
    For every vertex $v_{i}$ the clause $x_{i,1}\vee\cdots\vee x_{i,k}$.
  2. (2)
    For each pair of adjacent vertices $v_{i},v_{j}$ the clause $\bigwedge_{l=1}^{k}\neg x_{i,\ell}\vee\neg x_{j,\ell}$.
Now the chromatic number is the smallest $k$ for which the SAT problem above is satisfiable.
We used the MAGMA program chi-dimension-4.magma to perform the necessary computations and to generate the input for the SAT solver, for which we used PySAT [11]. This MAGMA program can be found as an ancillary file from the arXiv.org e-print archive. The MAGMA program contains the following:
  1. (1)
    The $52$ lattices associated to the $52$ inequivalent Delaunay subdivisions in dimension $4$. These are the lattices contained in Table 2.2 and Table 2.3;
  2. (2)
    A verification of Table 2.4: of the $52$ Cayley graphs associated to these lattices only $16$ are pairwise non-isomorphic;
  3. (3)
    The generation of SAT problems to confirm the Delaunay polytope bound and discrete torus bound. For each graph case we provide a number $c$ (which turns out to satisfy $c=\chi$) and
    1. (a)
      a certificate that $C_{1}$ is not $c-1$-colorable (DPB);
    2. (b)
      a certificate that the associated discrete torus $\Lambda/c\Lambda$ is $c$ colorable (DTB).
Here, we stick to the notation from Section 3 and Table 3.1.
4. Dimension 5 and onwards
We conclude the present discussion by some observations and open questions.
Firstly, the complete classification of inequivalent Delaunay subdivisions is also known in dimension $5$ by Dutour Sikirić, Garber, Schürmann, and Waldmann [7]. So our brute-force computational approach to find the chromatic number of all $5$-dimensional lattices is, in principle, applicable. However, all our computations for dimension $5$ remained inconclusive: The size of the finite graphs that need to be considered grows drastically, in any case where a SAT solver could provide an answer in reasonable time, the bounds were far apart. We think that strengthening the bounds is a more promising strategy than throwing a massive amount of computational time on the problem. For this reason we restricted to the $4$-dimensional case for now.
So, some questions remain:
  1. (1)
    Can we strengthen the bounds presented here, or find a way to compute them faster?
  2. (2)
    The maximal chromatic number in dimension 4 is 7, while the best known theoretical upper bound is $2^{4}$. Can one prove that indeed $2^{n-1}$ is a valid upper bound on the chromatic number of $n$-dimensional lattices for $n\geq 3$? It is true in dimension $3$ by the results in [8].
  3. (3)
    For all lattices studied in this work and in [8] we are aware of at least one periodic optimal coloring (while non-periodic optimal colorings may also exist). Now, given a general lattice, is it always true that there exists a periodic optimal coloring?
  4. (4)
    More generally, if $\Lambda^{\prime}$ is a sublattice of $\Lambda$, what can we say about $\Lambda^{\prime}$-periodic (optimal) colorings of $\Lambda$? Do they exist? If so, how many such? How small can the index $[\Lambda:\Lambda^{\prime}]$ become? Is there a bound $B(n)$ (depending on the dimension) such that there always exists a $\Lambda^{\prime}$-periodic optimal coloring with $[\Lambda:\Lambda^{\prime}]\leq B(n)$. In all known cases $[\Lambda:\Lambda^{\prime}]\leq\chi(\Lambda)^{n}$ is true.
Acknowledgments
The first named author likes to thank Mathieu Dutour Sikirić for helpful discussions.
The authors also express their gratitude to the suggestions of one of the anonymous referees. In comparison to an earlier draft of this manuscript we included a suggested strengthening of the statement of Lemmma 3.2, adapted question (2) from the referee report, and revised question (4) concerning periodic optimal colorings.

### v3

@@
 Equivalently, the chromatic number of a lattice $\Lambda$ is the chromatic number of the (infinite) Cayley graph on the additive group $\Lambda$ with strict Voronoi vectors as its set of generators.
-The strict Voronoi vectors of $\Lambda$, denoted by $\operatorname{Vor}(\Lambda)$, are the differences $v-w$ such that the intersection of the distinct translates $v+V(\Lambda)$ and $w+V(\Lambda)$ is a common facet.
+The Voronoi vectors of $\Lambda$, are the differences $v-w$ such that the intersection of the distinct translates $v+V(\Lambda)$ and $w+V(\Lambda)$ is a common face; but not necessarily a common facet.
+A Voronoi vector as above is a strict Voronoi vector if this common intersection is a facet, we denote the set of such vectors by $\operatorname{Vor}(\Lambda)$.
 Thus,
@@
 For instance, it is not known whether $\chi(\Lambda)$ is computable.
-By the compactness theorem of de Bruijn and Erdős [3] we only know that the decision problem “Is $\chi(\Lambda)\leq k$?” is semidecidable [^1] There is an algorithm which terminates with a positive answer after finitely many steps on every input $\Lambda$ with $\chi(\Lambda)\leq k$.
+By the compactness theorem of de Bruijn and Erdős [3] we only know that the decision problem “Is $\chi(\Lambda)\leq k$?” is semidecidable because $\chi(\Lambda)$ is equal to the largest chromatic number of all finite, induced subgraphs of $\Lambda$.
+This means that there is an algorithm which terminates with a positive answer after finitely many steps on every input $\Lambda$ with $\chi(\Lambda)\leq k$.
 On inputs $\Lambda$ with $\chi(\Lambda)>k$, the algorithm either terminates with a negative answer or it runs in an infinite loop.
-because $\chi(\Lambda)$ is equal to the largest chromatic number of all finite, induced subgraphs of $\Lambda$.
 Directly related is the question of whether one can always find a periodic coloring with $\chi(\Lambda)$ colors.
@@
 We prefer to work with Delaunay subdivisions, as they are more common in the literature.
-Note that our classification of Voroni graphs in dimension $4$ yields that there are $16$ such, up to equivalence, so their number is strictly smaller then the number of iso-edge domains.
+Note that our classification of Voroni graphs in dimension $4$ yields that there are $16$ such, up to equivalence, so their number is strictly smaller than the number of iso-edge domains.
 We now return to the discussion of Delaunay subdivisions. To state Voronoi’s result it is convenient to work with positive definite quadratic forms instead of lattices.
@@
 Here we recall some details of Voronoi’s classification of Delaunay triangulations of quaternary positive definite quadratic forms.
-We refer to the book by Schürmann [14] for a contemporary exhibition of Voronoi’s theory.
+We refer to the book by Schürmann [15] for a contemporary exposition of Voronoi’s theory.
 Let $\mathcal{D}$ be the Delaunay subdivision of some positive definite quadratic form $Q$, then
@@
 Under this group action there are only finitely many orbits of secondary cones.
-At the end of his second memoir, Voronoi [17] computed the orbits of full-dimensional secondary cones for $n\leq 4$. For $n\leq 3$ there is only one orbit and for $n=4$ there are exactly three orbits.
+At the end of his second memoir, Voronoi [18] computed the orbits of full-dimensional secondary cones for $n\leq 4$.
+For $n\leq 3$ there is only one orbit and for $n=4$ there are exactly three orbits.
+It turns out that in these cases, all full-dimensional secondary cones are simplicial.
+This is no longer the case for $n\geq 5$.
+For instance, in [7], for $n=6$, a full-dimensional secondary cone (of dimension $\binom{n(n+1)}{2}=21$) having $7145429$ many extreme rays was found.
 Using Voronoi’s notation, the three polyhedral domains ${\it\Delta}$, ${\it\Delta^{\prime}}$, ${\it\Delta^{\prime\prime}}$ are given by the extremal rays:
@@
 The column $r$ gives the regularity of the Voronoi graph, the column $|E|$ gives the number of edges, the column “order” gives the order of the automorphism group of the induced subgraph $\operatorname{Cayley}(\Lambda,\operatorname{Vor}(\Lambda))_{1}$.
-Here, we succeeded testing whether two Voronoi graphs are isomorphic in a rather ad hoc manner.
+Here we were able to test whether two Voronoi graphs are isomorphic in a rather ad hoc way.
 This can also be done in a systematic way: In the next section we provide a finite algorithm that can be used in general, for this we will briefly discuss how to test for isomorphism of general Cayley graphs associated with lattices with finite sets of generators; see also the paper [1] by Baburin which summarizes computational approaches to finding isomorphisms and automorphisms of Cayley graphs of lattices and contains references to earlier works in this field.
@@
 ∎
-For any choice of a full-dimenssional $\Lambda^{\prime}$ we refer to the bound of the above lemma as the discrete torus bound induced by $\Lambda^{\prime}$.
+For any choice of a full-dimensional $\Lambda^{\prime}$ we refer to the bound of the above lemma as the discrete torus bound induced by $\Lambda^{\prime}$.
 $(0,0)$$(1,0)$$(0,1)$$(1,1)$$(0,2)$$(2,0)$$(1,2)$$(2,1)$$(2,2)$
@@
 3.3. Proof of the main theorem
-The proof of our main theorem ultimately involves a finite list of computations.
+The proof of our main theorem ultimately involves a finite list of computations, although this was not guaranteed apriori.
 The general strategy is as follows.
@@
 We look for the smallest number $k$ such that there exists a coloring with $k$ color classes.
-This can be rephrased as a Boolean formula in conjunctive normal for (CNF) as follows.
+This can be rephrased as a Boolean formula in conjunctive normal form (CNF) as follows.
 Write $V=\{v_{1},\ldots,v_{n}\}$; these are the variables.
@@
 The authors also express their gratitude to the suggestions of one of the anonymous referees.
-In comparison to an earlier draft of this manuscript we included a suggested strengthening of the statement of Lemmma 3.2, adapted question (2) from the referee report, and revised question (4) concerning periodic optimal colorings.
+In comparison to an earlier draft of this manuscript we included a suggested strengthening of the statement of Lemma 3.2, adapted question (2) from the referee report, and revised question (4) concerning periodic optimal colorings.